# Sine laws on semigroups with an involutive anti-automorphism: A Levi–Civita approach via left translations

### v4

@@ Abstract.
which converts the problematic right translates into left translations. Using this left-translation approach, we obtain an anti-automorphic Levi–Civita closure principle and apply it to the generalized sine law, recovering the classical dichotomy $\beta\in\{\pm 1\}$ and the associated $xy$-addition law. Moreover, under a mild centrality assumption (e.g. for commutative semigroups), we also recover the standard $\sigma$-transformation rules (such as $f\circ\sigma=\beta f$ and the corresponding formula for $g\circ\sigma$).
@@
Functional equation, Sine addition law, Semigroup, Anti-automorphism, Levi-Civita equation, Central function.
@@ 1. Introduction
To explain the obstruction addressed in this paper, consider the two-parameter family of generalized sine addition/subtraction laws
@@
However, the situation changes when $\sigma$ is an anti-homomorphism, i.e. $\sigma(xy)=\sigma(y)\sigma(x)$. Then the natural right action no longer interacts with the product in the same way:
@@
This order reversal prevents a direct transfer of the usual Levi–Civita arguments based on right translates. Such anti-automorphisms arise naturally in many settings, for instance matrix transposition, inversion in groups, and involutions on symmetry structures. A principled semigroup theory in the anti-automorphic setting therefore calls for an adapted translation-operator framework.
Operator-level viewpoint. When $\sigma$ is an involutive anti-automorphism, the standard Levi–Civita method based on right translations breaks down because the family $\{R(\sigma(y))\}_{y\in S}$ satisfies the order-reversing composition law (1.2). We therefore work at the operator level: letting $J$ denote composition with $\sigma$, $(Jh)(x)=h(\sigma(x))$, we prove the conjugation identity
@@
which converts the problematic right translates into a left-translation action on the conjugated space $J(V)$.
Main idea: conjugation to left translations. Throughout, $\sigma:S\to S$ is assumed to be a surjective involutive anti-automorphism. The conjugation identity (1.3) (Lemma 3.1) allows us to transfer the right-translate action $R(\sigma(\cdot))$ to left translations, which restores the translation-invariance mechanism needed for a Levi–Civita analysis. In the application to the generalized sine law, a mild additional centrality hypothesis is only used to obtain the explicit $\sigma$-transformation rule for $g$.
Main results. Our first main result is an anti-automorphic Levi–Civita closure theorem (Theorem 3.2) for equations of the form
@@
with $\{f,g\}$ and $\{h_{1},h_{2}\}$ linearly independent. The theorem shows that $J(V)$ is invariant under left translations (equivalently, $V$ is invariant under $R(\sigma(\cdot))$), yielding a closure principle tailored to involutive anti-automorphisms.
Our second main result applies this closure to the generalized sine law (1.1). In Theorem 4.4 we obtain an analogue of Stetkær’s structural theorem: for linearly independent solutions $(f,g)$, one necessarily has $\beta\in\{\pm 1\}$ and the equation forces a simpler $xy$-addition law. Furthermore, assuming the existence of a central element $c\in Z(S)$ with $f(c)\neq 0$ (for instance, when $S$ is commutative and $f\not\equiv 0$), the solutions satisfy the expected $\sigma$-transformation rules:
@@
In particular, the anti-automorphic setting does not introduce new exceptional parameter regimes; instead, the conjugation viewpoint recovers the classical dichotomy.
Organization of the paper. Section 2 collects the needed definitions and notation (central/abelian functions, parity, and the regular actions $R$ and $L$). In Section 3 we develop the conjugation-based framework: we prove (1.3) and establish the anti-automorphic Levi–Civita closure theorem (Theorem 3.2). Section 4 applies this closure to the generalized sine law and proves our main structural classification (Theorem 4.4), where an additional mild centrality hypothesis is assumed only to obtain the explicit $\sigma$-transformation rule for $g$. Section 5 provides concrete illustrations from matrix groups and the symmetric group, where the conjugation identity can be seen explicitly. Finally, Section 6 gives concluding remarks and connections to recent centrality/abelianity results in the literature.
@@ 3. A Left-Translation Method via Operator Conjugation
Throughout this section, $\sigma:S\to S$ is a surjective anti-automorphism with $\sigma^{2}=\mathrm{id}$.
@@ 3.2. Anti-automorphism Levi–Civita closure
Let $\sigma:S\to S$ be a surjective involutive anti-automorphism. Suppose $f,g,h_{1},h_{2}:S\to F$ satisfy
@@
where $\{f,g\}$ and $\{h_{1},h_{2}\}$ are linearly independent. Set $V:=\mathrm{span}\{f,g\}$ and $V^{\sigma}:=J(V)=\mathrm{span}\{Jf,Jg\}$. Then, $V^{\sigma}$ is invariant under $L(y)$ for every $y\in S$. Equivalently, $V$ is invariant under $R(\sigma(y))$ for every $y\in S$.
@@
We now show that $L(y)(Jg)\in V^{\sigma}$ for all $y\in S$. Indeed, choose $y_{0}\in S$ such that $h_{2}(y_{0})\neq 0$ (exists since $h_{1},h_{2}$ are linearly independent). From
@@ 4. The Generalized Sine Law with an Anti-Automorphism
Lemma 4.3 (Central-element criterion for $\sigma$-conjugate span).
Let $S$ be a semigroup and let $\sigma:S\to S$ be a surjective involutive anti-automorphism. Assume that $f,u,h:S\to F$ satisfy
@@
If there exists a central element $c\in Z(S)$ such that $f(c)\neq 0$, then
@@
Multiplying by $\varepsilon$ and using $\varepsilon^{2}=1$, we get the identity of functions
@@
Now let $c\in Z(S)$ with $f(c)\neq 0$. Since $\sigma$ is surjective, pick $y_{0}\in S$ such that $\sigma(y_{0})=c$. Then
$(R(\sigma(y_{0}))f)(x)=f(x\sigma(y_{0}))=f(xc)=f(cx)=(L(c)f)(x),$
where we used $c\in Z(S)$. Applying Theorem 3.2 to (4.2) shows that $V:=\mathrm{span}\{f,u\}$ is invariant under $R(\sigma(\cdot))$, hence $R(\sigma(y_{0}))f\in V$, so $L(c)f\in V$.
Finally, take $y=c$ in (4.4):
$L(c)f=h(c)\,f+\varepsilon f(c)\,(u\circ\sigma).$
Since $L(c)f\in V$ and $f(c)\neq 0$, we solve for $u\circ\sigma$:
$u\circ\sigma=\frac{1}{\varepsilon f(c)}\,L(c)f-\frac{h(c)}{\varepsilon f(c)}\,f\in V.$
@@
Let $\sigma:S\to S$ be a surjective involutive anti-automorphism, and assume that there exists a central element $c\in Z(S)$ such that $f(c)\neq 0$. Suppose $f,g:S\to F$ are linearly independent and satisfy (4.1).
@@
From Step 1 we already have $f\circ\sigma=-f$, i.e. (4.3) holds with $\varepsilon=-1$. Applying Lemma 4.3 to (4.2), we conclude that
@@
$f(xy)=f(x)\bigl(g(\sigma(y))+\gamma f(\sigma(y))\bigr)-g(x)f(\sigma(y))=f(x)\bigl(-g(y)+\mu f(y)-\gamma f(y)\bigr)+g(x)f(y).$
@@
Comparing the two expressions for f(xy)-f(xy) yields $\gamma=0$. With $\gamma=0$, (4.15) becomes
@@
Applying $\sigma$ to $f(z)=\beta f(\sigma(z))$ and using $\sigma^{2}={\rm id}$ gives $\beta^{2}=1$. Since $\beta\neq-1$, we conclude
@@
This is exactly (4.2) with $u=h=g_{1}$ and $\varepsilon=1$ in (4.3). Applying Lemma 4.3, we get
@@
5. Examples: Concrete Illustrations of the Conjugation Identity
We illustrate the conjugation identity on familiar groups (hence semigroups). In each example, $\sigma$ is an involutive anti-automorphism and $J$ denotes composition with $\sigma$,
@@ 5. Examples: Concrete Illustrations of the Conjugation Identity
We emphasize that these examples are included solely to illustrate the conjugation identity $J\,R(\sigma(y))\,J=L(y)$ and the resulting left-translation action; they are independent of the additional centrality assumption used later in Theorem 4.4.
5.1. Example 1: $GL_{n}(K)$ with $\sigma(A)=A^{\top}$ and column-coordinate functions
Let $K$ be a field and let $S=GL_{n}(K)$ with $\sigma(A)=A^{\top}$. Then $\sigma$ is an involutive anti-automorphism since $(AB)^{\top}=B^{\top}A^{\top}$ and $(A^{\top})^{\top}=A$. Fix $v\in K^{n}$ with $v\neq 0$ and define, for $i=1,\dots,n$,
@@ 5.1. Example 1: $GL_{n}(K)$ with $\sigma(A)=A^{\top}$ and column-coordinate functions
5.2. Example 2: The symmetric group $S_{n}$ with $\sigma(\pi)=\pi^{-1}$
@@
5.3. Example 3: $SO(3,\mathbb{R})$ with $\sigma(g)=g^{-1}=g^{\top}$ and column functions
Let $S=SO(3,\mathbb{R})$ and $\sigma(g)=g^{-1}=g^{\top}$. Write $g=(g_{ij})_{1\leq i,j\leq 3}$ and define
$f_{1}(g):=g_{11},\qquad f_{2}(g):=g_{21},\qquad f_{3}(g):=g_{31},\qquad V:=\mathrm{span}\{f_{1},f_{2},f_{3}\}.$
Verification of $JR(\sigma(y))J=L(y)$. We have $(Jf_{i})(X)=f_{i}(X^{\top})=X_{1i}$. For $y\in SO(3,\mathbb{R})$ and $x\in SO(3,\mathbb{R})$,

### v5

@@ Abstract.
which converts the problematic right translates into left translations. Using this left-translation approach, we obtain an anti-automorphic Levi–Civita closure principle and apply it to the generalized sine law. Remarkably, the classical dichotomy $\beta\in\{\pm 1\}$ and the parity relation $f\circ\sigma=\beta f$ are recovered unconditionally. Furthermore, under a natural bridge hypothesis, which is automatically satisfied when there exists a central element $c$ with $f(c)\neq 0$, we obtain the corresponding standard $xy$-addition law and the exact $\sigma$-transformation rule for $g$.
@@
Functional equation, Sine addition law, Semigroup, Anti-automorphism, Levi–Civita equation, Translation invariance.
@@ 1. Introduction
To motivate the operator viewpoint adopted here, consider the two-parameter family of generalized sine addition/subtraction laws
@@
When $\sigma$ is an anti-homomorphism, i.e. $\sigma(xy)=\sigma(y)\sigma(x)$, the parametrized family $y\mapsto R(\sigma(y))$ satisfies the order-reversing composition law
@@
Thus the family $R(\sigma(\cdot))$ is no longer a right-regular representation in the parameter $y$. Of course, because $\sigma^{2}=\mathrm{id}$, one may rewrite (1.1) by replacing $y$ with $\sigma(y)$ and thereby recover a standard $xy$-equation. Our aim here is complementary: we isolate an intrinsic operator identity explaining how the anti-automorphic family $R(\sigma(\cdot))$ is converted into a genuine translation action.
Operator-level viewpoint. When $\sigma$ is an involutive anti-automorphism, the family $\{R(\sigma(y))\}_{y\in S}$ is order-reversing in the parameter $y$ by (1.2). Letting $J$ denote composition with $\sigma$, $(Jh)(x)=h(\sigma(x))$, we prove the conjugation identity
@@
which converts the family $R(\sigma(\cdot))$ into left translations. This gives a natural operator-level reformulation of the anti-automorphic situation and is the viewpoint used throughout the paper.
Main idea: conjugation to left translations. Throughout, $\sigma:S\to S$ is assumed to be an involutive anti-automorphism. The conjugation identity (1.3) (Lemma 3.1) allows us to transfer the right-translate action $R(\sigma(\cdot))$ to left translations, which restores the translation-invariance mechanism needed for a Levi–Civita analysis. In the application to the generalized sine law, the decisive additional input is a bridge condition of the form
$\exists\,d\in S:\qquad f(d)\neq 0\quad\text{and}\quad L(d)f\in\mathrm{span}\{f,g\}.$
A central element $c\in Z(S)$ with $f(c)\neq 0$ is a convenient sufficient condition for this bridge assumption, but the bridge formulation is the intrinsic one used in the proof.
Main results. Our first main result, Theorem 3.2, is an anti-automorphic Levi–Civita closure theorem for equations of the form
@@
where $\{f,g\}$ is linearly independent and $h_{2}\not\equiv 0$. The theorem shows that $J(V)$ is invariant under left translations (equivalently, $V$ is invariant under $R(\sigma(\cdot))$), yielding a closure principle tailored to involutive anti-automorphisms.
Our second main result applies this closure to the generalized sine law (1.1). In Theorem 4.6, under the bridge assumption above, we obtain an analogue of Stetkær’s structural theorem: for linearly independent solutions $(f,g)$, one necessarily has $\beta\in\{\pm 1\}$, the equation yields the standard $xy$-addition law, and the expected $\sigma$-transformation rules
@@
Moreover, the proof shows that the parity relation $f\circ\sigma=\beta f$ is obtained before the bridge assumption is invoked, whereas the bridge assumption is used only to obtain control of $g\circ\sigma$ and thereby to reach the standard $xy$-addition law. Corollary 4.7 records the frequently useful central-element criterion, and Example 5.4 shows that the bridge assumption is strictly weaker.
Organization of the paper. Section 2 collects the needed definitions and notation (central/abelian functions, parity, and the regular actions $R$ and $L$). In Section 3 we develop the conjugation-based framework: we prove (1.3) and establish the anti-automorphic Levi–Civita closure theorem (Theorem 3.2). Section 4 applies this closure to the generalized sine law and proves our main structural classification under the bridge assumption; Corollary 4.7 gives the corresponding result for central elements. Section 5 provides concrete illustrations from matrix groups and the symmetric group, including a non-central example showing that the bridge assumption can hold even when no central element $c$ satisfies $f(c)\neq 0$. Finally, Section 6 gives concluding remarks and connections to recent centrality/abelianity results in the literature.
@@ 3. A Left-Translation Method via Operator Conjugation
Throughout this section, $\sigma:S\to S$ is an involutive anti-automorphism.
@@ 3.2. Anti-automorphism Levi–Civita closure
Let $\sigma:S\to S$ be an involutive anti-automorphism. Suppose $f,g,h_{1},h_{2}:S\to F$ satisfy
@@
where $\{f,g\}$ is linearly independent and $h_{2}\not\equiv 0$. Set $V:=\mathrm{span}\{f,g\}$ and $V^{\sigma}:=J(V)=\mathrm{span}\{Jf,Jg\}$. Then $V^{\sigma}$ is invariant under $L(y)$ for every $y\in S$. Equivalently, $V$ is invariant under $R(\sigma(y))$ for every $y\in S$.
@@
We now show that $L(y)(Jg)\in V^{\sigma}$ for all $y\in S$. Indeed, choose $y_{0}\in S$ such that $h_{2}(y_{0})\neq 0$ (this exists because $h_{2}\not\equiv 0$). From
@@
Remark 3.3 (Sharpness of the hypothesis $h_{2}\not\equiv 0$).
Theorem 3.2 does not require linear independence of $h_{1}$ and $h_{2}$; in particular, it applies when $h_{1}=0$ and $h_{2}\not\equiv 0$. The proof uses only the existence of a point $y_{0}\in S$ with $h_{2}(y_{0})\neq 0$.
This hypothesis is sharp. If $h_{2}\equiv 0$, the conclusion can fail. For example, let $S=C_{3}=\{e,a,a^{2}\}$, let $\sigma=\mathrm{id}_{S}$, let $f\equiv 1$, $h_{1}\equiv 1$, $h_{2}\equiv 0$, and let $g=\mathbf{1}_{\{e\}}$. Then
$f(x\sigma(y))=1=f(x)h_{1}(y)+g(x)h_{2}(y)\qquad(\forall x,y\in S),$
and $f,g$ are linearly independent. However,
$L(a)g=\mathbf{1}_{\{a^{2}\}}\notin\mathrm{span}\{1,\mathbf{1}_{\{e\}}\}=\mathrm{span}\{f,g\}.$
Hence $V^{\sigma}=V$ is not invariant under left translations.
@@ 4. The Generalized Sine Law with an Anti-Automorphism
Lemma 4.3 (Bridge criterion for $\sigma$-conjugate span).
Let $S$ be a semigroup and let $\sigma:S\to S$ be an involutive anti-automorphism. Assume that $f,u,h:S\to F$ satisfy
@@
If there exists $d\in S$ such that
$f(d)\neq 0\qquad\text{and}\qquad L(d)f\in\mathrm{span}\{f,u\},$ (4.4)
then
@@
Multiplying by $\varepsilon$ and using $\varepsilon^{2}=1$, we get
@@
Set $V:=\mathrm{span}\{f,u\}$. Taking $y=d$ in (4.5) and using (4.4), we obtain
$L(d)f=h(d)\,f+\varepsilon f(d)\,(u\circ\sigma)\in V.$
Since $f(d)\neq 0$, we can solve for $u\circ\sigma$:
$u\circ\sigma=\frac{1}{\varepsilon f(d)}\bigl(L(d)f-h(d)f\bigr)\in V.$
@@
Corollary 4.4 (Central-element criterion).
Let $S$ be a semigroup and let $\sigma:S\to S$ be an involutive anti-automorphism. Assume that $f,u,h:S\to F$ satisfy (4.2), where $f$ and $u$ are linearly independent, and that (4.3) holds for some $\varepsilon\in\{\pm 1\}$. If there exists a central element $c\in Z(S)$ such that $f(c)\neq 0$, then
$u\circ\sigma\in\mathrm{span}\{f,u\}.$
Proof.
Set $V:=\mathrm{span}\{f,u\}$ and let $y_{0}:=\sigma(c)$. Then $\sigma(y_{0})=c$, so by (4.2),
$R(c)f=R(\sigma(y_{0}))f=h(y_{0})\,f+f(y_{0})\,u\in V.$
Since $c\in Z(S)$, we have $L(c)f=R(c)f\in V$. Thus (4.4) holds with $d=c$, and Lemma 4.3 gives the desired conclusion.
For the commutative case, choose $d\in S$ with $f(d)\neq 0$. Since $\sigma$ is bijective, $d=\sigma(\sigma(d))$, and (4.2) with $y=\sigma(d)$ gives
$R(d)f=R(\sigma(\sigma(d)))f\in V.$
Because $S$ is commutative, $L(d)f=R(d)f\in V$, so (4.4) holds and Lemma 4.3 applies. ∎
Remark 4.5.
Strictly speaking, the algebraic derivations in Theorem 3.2, Lemma 4.3, and Corollary 4.4 do not rely on the linear independence of the functions involved; their conclusions remain valid even if the function pairs are linearly dependent. We state them with the linear independence hypothesis primarily for contextual consistency, as it seamlessly sets up the non-degenerate framework required for our main structural classification in the following important theorem.
@@
Let $\sigma:S\to S$ be an involutive anti-automorphism. Suppose $f,g:S\to F$ are linearly independent and satisfy (4.1). Assume that there exists $d\in S$ such that
$f(d)\neq 0\qquad\text{and}\qquad L(d)f\in\mathrm{span}\{f,g\}.$ (4.6)
@@
From Step 1 we already have $f\circ\sigma=-f$, i.e. (4.3) holds with $\varepsilon=-1$. Since u=gu=-g, the bridge hypothesis (4.6) reads
$f(d)\neq 0\qquad\text{and}\qquad L(d)f\in\mathrm{span}\{f,u\}.$
Applying Lemma 4.3 to (4.2), we conclude that
@@
$f(xy)=f(x)\bigl(g(\sigma(y))+\gamma f(\sigma(y))\bigr)-g(x)f(\sigma(y))=f(x)\bigl(-g(y)+(\mu-\gamma)f(y)\bigr)+g(x)f(y).$
@@
Comparing the two expressions for f(xy)-f(xy), we obtain
$2\gamma\,f(x)f(y)=0\qquad(\forall x,y\in S).$
Since $f\not\equiv 0$, choose $x_{0}\in S$ with $f(x_{0})\neq 0$; taking $x=y=x_{0}$ and using $\operatorname{char}(F)\neq 2$, we get $\gamma=0$.
With $\gamma=0$, (4.17) becomes
@@
Applying $\sigma$ to $f(z)=\beta f(\sigma(z))$ and using $\sigma^{2}={\rm id}$, we obtain
$f(\sigma(z))=\beta f(z)\qquad(\forall z\in S).$
Hence
$f(z)=\beta f(\sigma(z))=\beta^{2}f(z)\qquad(\forall z\in S).$
Thus $(1-\beta^{2})f(z)=0$ for all $z\in S$. Since $f\not\equiv 0$, choose $z_{0}\in S$ with $f(z_{0})\neq 0$. Then $(1-\beta^{2})f(z_{0})=0$, hence $\beta^{2}=1$. Since $\beta\neq-1$, we conclude
@@
This is exactly (4.2) with $u=h=g_{1}$ and $\varepsilon=1$ in (4.3). Since $\mathrm{span}\{f,g_{1}\}=\mathrm{span}\{f,g\}$, the bridge hypothesis (4.6) gives
$f(d)\neq 0\qquad\text{and}\qquad L(d)f\in\mathrm{span}\{f,g_{1}\}.$
Applying Lemma 4.3, we get
@@
Corollary 4.7.
The conclusions of Theorem 4.6 hold in particular if there exists a central element $c\in Z(S)$ such that $f(c)\neq 0$.
Proof.
By (4.1), taking $y=\sigma(c)$ shows that
$R(c)f=R(\sigma(\sigma(c)))f\in\mathrm{span}\{f,g\}.$
Since $c\in Z(S)$, we have $L(c)f=R(c)f\in\mathrm{span}\{f,g\}$. Hence (4.6) holds with $d=c$, and Theorem 4.6 applies. ∎
Remark 4.8.
In the proof of Theorem 4.6, the dichotomy $\beta\in\{\pm 1\}$ and the parity relation $f\circ\sigma=\beta f$ are obtained before the bridge hypothesis (4.6) is used. The bridge assumption enters only through Lemma 4.3, which is invoked to show that $g\circ\sigma$ (or $g_{1}\circ\sigma$ in the case $\beta\neq-1$) belongs to the span of the basic solution pair. This span-closure step is then used to derive the standard $xy$-addition law and the explicit $\sigma$-transformation rule for $g$. The central-element condition of Corollary 4.7 is merely a convenient sufficient condition for (4.6); Example 5.4 in the sequel shows that it is strictly stronger.
5. Examples and Further Illustrations
We first illustrate the conjugation identity on familiar groups (hence semigroups). In each of the first three examples, $\sigma$ is an involutive anti-automorphism and $J$ denotes composition with $\sigma$,
@@ 5. Examples and Further Illustrations
We emphasize that Examples 1–3 below are included solely to illustrate the conjugation identity $J\,R(\sigma(y))\,J=L(y)$ and the resulting left-translation action (independent of the bridge hypothesis used in Theorem 4.6), whereas Example 4 focuses on this bridge hypothesis to show that it may hold even when no central element $c$ satisfies $f(c)\neq 0$.
5.1. Example 1: GL(n,K) with transposition and column-coordinate functions
Let $K$ be a field and let $S=GL_{n}(K)$ with $\sigma(A)=A^{\top}$. In this example, we take the coefficient field to be $F=K$. Then $\sigma$ is an involutive anti-automorphism since $(AB)^{\top}=B^{\top}A^{\top}$ and $(A^{\top})^{\top}=A$. Fix $v\in K^{n}$ with $v\neq 0$ and define, for $i=1,\dots,n$,
@@ 5.1. Example 1: GL(n,K) with transposition and column-coordinate functions
5.2. Example 2: The symmetric group Sn
@@
5.3. Example 3: SO(3,R) and column functions
Let $S=SO(3,\mathbb{R})$ and $\sigma(g)=g^{-1}=g^{\top}$. Here the coefficient field is $F=\mathbb{R}$. Write $g=(g_{i,\,j})_{1\leq i,\,j\leq 3}$ and define
$f_{1}(g):=g_{1,\,1},\qquad f_{2}(g):=g_{2,\,1},\qquad f_{3}(g):=g_{3,\,1},\qquad V:=\mathrm{span}\{f_{1},f_{2},f_{3}\}.$
Verification of $JR(\sigma(y))J=L(y)$. We have $(Jf_{i})(X)=f_{i}(X^{\top})=(X^{\top})_{i,\,1}=X_{1,\,i}$. For $y\in SO(3,\mathbb{R})$ and $x\in SO(3,\mathbb{R})$,